### v1

Abstract
To solve ever more complex problems, Deep Neural Networks are scaled to billions of parameters, leading to huge computational costs. An effective approach to reduce computational requirements and increase efficiency is to prune unnecessary components of these often over-parameterized networks. Previous work has shown that attribution methods from the field of eXplainable AI serve as effective means to extract and prune the least relevant network components in a few-shot fashion. We extend the current state by proposing to explicitly optimize hyperparameters of attribution methods for the task of pruning, and further include transformer-based networks in our analysis. Our approach yields higher model compression rates of large transformer- and convolutional architectures (VGG, ResNet, ViT) compared to previous works, while still attaining high performance on ImageNet classification tasks. Here, our experiments indicate that transformers have a higher degree of over-parameterization compared to convolutional neural networks. Code is available at https://github.com/erfanhatefi/Pruning-by-eXplaining-in-PyTorch.
Keywords:
Explainable AI Pruning Attribution Optimization
1 Introduction
Figure 1: We propose a pruning framework based on optimizing attribution methods from the field of XAI. Compared to random pruning, pruning the least relevant structures first (“relevant” according to an XAI attribution method of choice, and indicated by red color), has been shown to result in an improved performance-sparsity tradeoff (simplified illustration depicted). By optimizing attribution methods specifically for pruning, we can reduce the tradeoff even further.
In recent years, Deep Neural Networks (DNNs) have been growing increasingly larger, demanding more computational resources and memory. To address these challenges, several efficient architectures, such as MobileNet [23] or EfficientFormer [29], have been proposed to reduce computational costs. However, the gain in efficiency comes at the cost of performance, as inherently efficient architectures struggle to keep pace with the recent surge in high-performing transformer models.
One widely adopted approach within the community of efficient Deep Learning is quantization [18, 50], which involves compressing model parameters or features by reducing the number of bits used for representation, potentially followed by re-training to regain lost model performance. Another effective approach is model sparsification, commonly referred to as pruning, where irrelevant structures within the model are removed to reduce complexity and improve efficiency. It is to note, that quantization and pruning can be applied in tandem [26]. In this work, unlike previous studies that mandate post-retraining in their framework [15, 36], we focus on pruning without additional training.
The key challenge in pruning (without sequential re-training) is the identification of structures within the model that can be removed without adversely affecting its performance. Earlier works have shown that structures may be chosen according to, e.g., parameter magnitudes [19, 28], functional redundancy [17], sensitivity [48, 32], or importance in the decision-making process [49, 7]. Especially the latter approach has been promising, utilizing tools from the seemingly unrelated field of XAI that originally seeks to explain the model reasoning process to human stakeholders.
Backpropagation-based attribution methods, particularly Layer-wise Relevance Propagation (LRP) [5, 33, 34], play a crucial role in this context. LRP works by tracing the internal reasoning of the model, assigning contribution scores to latent neurons, which highlight their importance to the final prediction. These scores provide a quantitative basis for identifying (ir-)relevant structures and subgraphs within the model that are (un-)essential for making accurate predictions, thus guiding the pruning process.
While prior attribution-based pruning [49] achieved remarkable results, it is restricted to heuristically chosen LRP hyperparameters and limited to specific Convolutional Neural Networks (CNNs) and toy models.
Following up upon the work [35] that optimized LRP w.r.t. specific explainability metrics, we argue that LRP or any other attribution method can be optimized specifically for the goal of more effective pruning (see Fig. 1). In addition, recent works [2, 4] have extended LRP to transformer-based architectures, allowing us now to also apply attribution-based pruning to the Vision Transformer (ViT) architecture.
In this work, we revisit attribution-based model sparsification, and extend the current state-of-the-art by
  1. 1.
    proposing a novel pruning framework based on optimizing attribution method hyperparameters to achieve higher sparsification rates.
  2. 2.
    incorporating transformer-based architectures, such as ViTs, to our framework by using recently developed attribution methods.
  3. 3.
    discussing differences between CNNs and ViTs in terms of pruning and over-parameterization.
  4. 4.
    revealing that attributions optimized for explanations are not necessarily the best for pruning. Notably, simple hyperparameter settings show high performance across all tested architectures.
2 Related Work
In the following, we introduce related works in the field of XAI, efficient Deep Learning and the intersection between both.
Explainability and Local Feature Attribution Methods
Research in local explainability led to a plethora of methods [30, 37], commonly resulting in local feature attributions quantifying the importance of input features in the decision-making process. These attributions are often shown in the form of heatmaps in the vision domain. Notably, methods based on (modified) gradients, backpropagate attributions from the output to the input through the network, conveniently offering attributions of all latent components and neurons in a single backward pass [44, 42]. Gradient-based attribution methods, however, can suffer from noisy gradients, rendering them unreliable for deep architectures [6]. Prominently, LRP [5, 33] introduces a set of different rules (with hyperparameters) that allow to reduce the noise level. In fact, as shown in [35], attribution methods such as LRP can be optimized for certain XAI criteria, e.g., faithfulness or complexity [33, 22]. We follow up on this observation, and specifically optimize XAI w.r.t. pruning.
Pruning of Deep Neural Networks
For pruning CNNs, either individual (kernel) weights or whole structures, e.g., filters of convolution layers or neurons can be recognized as candidates for pruning [21]. For transformer architectures, such structures include heads of attention modules or linear layers inside transformer blocks [46, 27]. In order to prune such structures, several criteria have been proposed to indicate which components are best suited to be removed, retaining performance as best as possible. The work of [19] suggests pruning parameters based on weight magnitudes, offering a computationally-free criterion. Alternatively, in [11], the authors propose to prune neurons based on their activation patterns. However, recent work [7] has shown that weight or activation magnitudes do not necessarily reflect a component’s contribution during the inference process, e.g., also a small weight has the potential to be very relevant for the predication of a class.
Explainability For Efficient Deep Learning
The work of [49] introduces a novel pruning criterion based on XAI, and proposes to use latent relevance values extracted from the attribution method of LRP [5, 34] to assign importance scores to network structures. By taking into account how structures are used during the inference process, [49] can effectively improve the efficiency-performance tradeoff. Later, the works of [7, 43] also illustrate the value of XAI methods for network quantization. In all these works, heuristically chosen hyperparameters for LRP have been used, which overlooks the potential for specific optimization to the task of pruning. Notably, LRP is model-specific and thus restricted to compatible architectures. In this work, we also include recent LRP extensions to transformer architectures, and observe, that simple LRP settings commonly result in high pruning performances.
3 Methods
This work proposes a framework for pruning DNNs using attribution methods from the field of XAI with hyperparameters specifically optimized for sparsification. We begin with presenting our method in the form of a general XAI-based pruning principle in Sec. 3.1, followed by introducing LRP attributions and corresponding hyperparameters suitable for optimization, in Sec. 3.2 and Sec. 3.3, respectively. Lastly, Sec. 3.4 describes our optimization methodology.
3.1 Attribution-based Pruning
For our structured pruning framework, we view a DNN as a collection of $p$ (interlinked) components $\Psi=\{\psi_{1},\dots,\psi_{p}\}$, that can correspond to, e.g., whole layers, (groups of) neurons, convolutional filters or attention heads. We further assume access to an attribution method that generates attribution scores (relevance) $R_{\psi_{k}}({x_{i}})$ of component $\psi_{k}\in\Psi$ for the prediction of a sample $x_{i}$. The overall relevance of $\psi_{k}$ is then estimated through the mean relevance over a set of reference samples $\mathcal{X}_{\text{ref}}=\{x_{1},x_{2},\dots,x_{n_{\text{{ref}}}}\}$ as
Figure 2: Attribution-based pruning workflow: Firstly, the relevant model structures are found by explaining a set of reference samples. The attribution method of choice (here LRP) highlights the components and paths in the network which positively and negatively contribute to the decision-making. Positive and negative relevances are indicated by red and blue color respectively, and components with white color indicate no relevance. Removing structures that receive the least relevance results in a sparser subnetwork, which performs significantly better than after random pruning. Notably, relevances can be computed w.r.t. a subset of output classes (e.g., “corgi” only), resulting in a subnetwork specifically designed to perform the restricted task. Credit: Nataba/iStock.
$\bar{R}_{\psi_{k}}=\frac{1}{n_{\text{ref}}}\sum_{i=1}^{n_{\text{ref}}}R_{\psi_{k}}({x_{i}})\,.$ (1)
Ultimately, we collect relevances for all components via the set $\mathcal{R}$, given as
=$\mathcal{R}=\{\bar{R}_{\psi_{1}},\bar{R}_{\psi_{2}},\dots,\bar{R}_{\psi_{p}}\}\,,$ (2)
which, in turn, allows to define a pruning order for the model components. Specifically, the indices $c$ for the components to be pruned up to the $q$-th place are given by
$\centering\{c\}_{q}=\text{argsort}(\mathcal{R})_{1,2,\dots,q}\,\@add@centering$ (3)
resulting in the set of least relevant components. Finally, the $q$ least relevant components are pruned by removing or masking the components from the computational graph as
$\centering\forall{\psi_{i}\in\Psi}:\psi_{i}\mapsto(1-\textbf{1}_{i\in\{c\}_{q}})\psi_{i}\@add@centering$ (4)
where 1 is an indicator function with condition $i\in\{c\}_{q}$. The whole attribution-based pruning workflow is depicted in Fig. 2. Notably, previous work [49] demonstrates that LRP is an effective method for attributing latent structures, as used for Eq. 1. LRP further offers several hyperparameters to tune, which makes it a versatile choice for our framework.
3.2 Layer-wise Relevance Propagation
Layer-wise Relevance Propagation [5, 34] is a rule-based backpropagation algorithm that was designed as a tool for interpreting non-linear learning models by assigning attribution scores, called “relevances”, to network units proportionally to their contribution to the final prediction value. Unlike other gradient- or perturbation-based methods, LRP treats a neural network as a layered directed acyclic graph with $L$ layers and input $x$:
$\centering f(x)=f^{L}\circ\dots\circ f^{l}\circ f^{l-1}\circ\dots\circ f^{1}(x)\@add@centering$ (5)
Beginning with an initial relevance score $R_{j}^{L}$ at output $j$ of layer $f^{L}$, the score is layer-by-layer redistributed to its input variables depending on the contribution from the input variables to the output value.
Given a layer, we consider its pre-activations $z_{ij}$ mapping inputs $i$ to outputs $j$ and their aggregations $z_{j}=\sum_{i}z_{ij}$. Commonly in linear layers such a computation is given with $z_{ij}=a_{i}w_{ij}$, where $w_{ij}$ are its weight parameters and $a_{i}$ the activation of neuron $i$.
Then, LRP distributes relevance quantities $R_{j}^{l}$ received from upper layers towards lower layers proportionally to the relative contributions of $z_{ij}$ to $z_{j}$, i.e.,
$\centering R_{i\leftarrow j}^{(l-1,l)}=\frac{z_{ij}}{z_{j}}R_{j}^{l}\,.\@add@centering$ (6)
In other words, the relevance message $R_{i\leftarrow j}^{(l-1,l)}$ quantifies the contribution of neuron $i$ at layer $l-1$, to the activation of neuron $j$ at layer $l$.
To obtain the contribution of neuron $i$ to all upper layer neurons $j$, all incoming relevance messages $R_{i\leftarrow j}^{(l-1,l)}$ are losslessly aggregated as
$R_{i}^{l-1}=\sum_{j}R_{i\leftarrow j}^{(l-1,l)}.$ (7)
This process ensures the property of relevance conservation between adjacent layers:
$\sum_{i}R^{l-1}_{i}=\sum_{i,j}R^{(l-1,l)}_{i\leftarrow j}=\sum_{j}R^{l}_{j}$ (8)
which guarantees that the sum of all relevance in each layer stays the same.
When a group of neurons performs the same task within a convolutional channel or attention head, it is beneficial to aggregate the total relevance of the entire group into a single relevance score. This aggregation process helps in simplifying the analysis and interpretation of the model’s behavior by focusing on the collective relevance of each convolutional channel or attention head, rather than examining individual neurons. Further discussions of component-wise aggregation are given in Appendix 0.B.
3.3 Tuneable Hyperparameters of LRP
Within the LRP framework, various rules (as detailed in Sec. 0.A.2) have been proposed to tune relevance computation for higher explainability.
Further, several works have shown that using different rules for different architecture parts (referred to as a “composite”) [33, 25, 2] is favorable for explanations as it improves robustness and faithfulness. We follow up on this idea, and perform a large-scale analysis over a variety of rules for different network parts.
Inspired by the tuning settings of [35], we split a network into four parts:
  1. 1.
    the final classification layers, denoted as Fully-Connected Layers (FCL),
  2. 2.
    the last 25% of hidden layers before the FCL, denoted as High-Level hidden Layers (HLL),
  3. 3.
    the first 25% of hidden layers, denoted as Low-Level hidden Layers (LLL),
  4. 4.
    and the remaining hidden layers between HLL and LLL, denoted as Mid-Level hidden Layers (MLL)
For each group, a specific rule can be chosen, which together ultimately form the LRP composite. We refrain from treating each individual layer as a separate group due to the increased computational expenses associated with this setting.
For transformer architectures, the works of [2, 4] proposed novel LRP rules to attribute softmax non-linearities in the attention modules (see Sec. 0.A.2 for more details), encouraging us to also optimize the LRP rule for the softmax non-linearities, resulting in an additional hyperparameter here.
It is important to note, that relevance values can be positive or negative. Components with positive relevance values indicate that they significantly contribute towards the class prediction. Conversely, components with negative relevance values diminish the classification score, effectively speaking against the predicted class, as visible in Fig. 2 where “cat” features (and the respective model components) speak against the “corgi” prediction. Components with near zero relevance do not contribute to the inference process in any meaningful way. Thus, a question arises w.r.t. the optimal order of pruning: Should we start with negative, or near zero relevant components? To that end, we also add a hyperparameter indicating whether absolute relevance values are used as in Eq. 1, i.e., $\bar{R}_{\psi_{k}}\mapsto|\bar{R}_{\psi_{k}}|$ for all $\psi_{k}\in\Psi$ if true. The work of [49] takes into account only positive attributions by their LRP rule choice (see LRP-$z^{+}$ in Sec. 0.A.2). Our experiments in Appendix 0.F, however, reveal that the highest pruning rates are achieved by first pruning components with near zero relevance, as their minimal contribution ensures low impact on the overall model performance.
3.4 Hyperparameter Optimization Procedure
In order to optimize the hyperparameters of an attribution method, we first define an optimization objective $C$. As we perform our analysis for classification tasks, we measure the top-1 accuracy on the validation dataset. In principle, a different performance criterion can be chosen here. Concretely, we measure model performance for different pruning rates $\text{PR}_{i}=\frac{1}{m}i$ as given after step $i\in\{0,\dots m-1\}$ of in total $m$ steps (excluding a 100% rate). After sequentially increasing the pruning rate, and plotting performance against sparsity, we receive a curve that indicates the sparsity-accuracy tradeoff as, e.g., shown in Fig. 1 (left). We therefore propose to measure the resulting area under the curve $A_{\text{PR}}$ as given by
$\centering A_{\text{PR}}=\frac{1}{m}\sum_{i=0}^{m-1}C(f_{\Psi_{\text{PR}_{i}}(\theta)})\@add@centering$ (9)
where the performance as given by $C$ depends on the network $f$ and its pruned parameters $\Psi_{\text{PR}_{i}}(\theta)$ after pruning step $i$ using the hyperparameter setting $\theta$ for the attribution method. Ultimately, our objective will be the maximization of $A_{\text{PR}}$ w.r.t. $\theta$. Hyperparameter search is performed via grid search and Bayesian optimization techniques, as detailed in Appendix 0.D.
4 Experiments
We begin our experiments with exploring the over-parameterization problem of DNNs in Sec. 4.1. This is followed by our results on finding the best LRP hyperparameters for pruning CNNs and ViTs in Sec. 4.2 and Sec. 4.3, respectively. Lastly, we compare the effect of pruning using ideal and random attributions in Sec. 4.4 by evaluating how explanation heatmaps change after pruning.
Experimental Setting
In our experiments, we optimize LRP hyperparameters for pruning convolution filters of VGG-16 [41] (with and without BatchNorm [24] layers), ResNet-18 and ResNet-50 [20] architectures (with 4224, 4224, 4800 and 26560 filters overall, respectively), as well as linear layers and attention heads of the ViT-B-16 transformer [12] (with 46080 neurons and 144 heads). All models are pre-trained [31] and evaluated on the ImageNet dataset [10]. For hyperparameter optimization, we measure model performance for 20 pruning rates (from 0 % up to 95 %) on the validation dataset. To compute latent attributions, a set of reference samples has been chosen from the training set of ImageNet (different set sizes are discussed later in Sec. 4.2 and 4.3).
4.1 How Over-Parameterized are Vision Models?
Training a large DNN from scratch to solve a specific task can be computationally expensive. A popular approach for saving training resources is to instead fine-tune a large pre-trained (foundation) model, which often requires fewer training epochs and provides high (or even higher) model generalization, especially in sparse data settings [9]. Notably, one of the effects that arise when solving for simple(r) tasks, is that we likely end up with an over-parameterized model as only a subset of very specialized latent features are necessary to solve the task, which makes pruning especially interesting in this case.
Figure 3: Investigating over-parameterization in DNNs through attribution-based pruning (LRP-$\epsilon$ in blue, unoptimized in orange color) and random pruning (green color). We compare pruning of all models w.r.t. different task difficulties, i.e., to differentiate between 1000 (dotted line), 100 (dashed line) or three ImageNet classes (solid line). High performance for high sparsification rates indicates over-parameterization, i.e., many network components are not important for the task. Compared to the ResNet-18 and VGG-16-BN CNNs (top), the ViT-B-16 transformer shows a higher degree of over-parameterization (bottom). SEM is illustrated (shaded area) in the current and all other figures.
In fact, when pruning the ResNet-18 and VGG-16-BN models that were pre-trained to detect all of the 1000 ImageNet classes, parameter count can only be reduced by a couple of percent without meaningful degradation of model performance, as shown in Fig. 3 (top). This indicates that ImageNet itself is a complex task, and most of the trained parameters are actually relevant. However, as the task becomes simpler (simulated by reducing the number of output classes to 100 or three in the evaluation), pruning rates can be increased to a much higher level without critical accuracy loss.
For baseline comparisons, we utilize random pruning, Integrated Gradient [44] and our optimized LRP attributions (see Sec. 4.2). We intentionally exclude naive weight magnitude or activation pruning since [49] has previously demonstrated that XAI-based methods outperform these traditional approaches. Notably, especially for simpler tasks, attribution-based pruning outperforms random pruning, as attributions are output-specific and allow identifying the sub-graph that is relevant for the restricted task, as also illustrated in Fig. 2.
Compared to the previously discussed CNNs, the ViT-B-16 vision transformer seems to be more over-parameterized, as visible in Fig. 3 (bottom). Up to a pruning rate of 20 %, no meaningful accuracy loss is measured for detecting the 1000 ImageNet classes, irrespective of whether attention heads or neurons in linear layers are pruned. This might be due to the fact, that the vision transformer consists of about 46 thousand neurons, which is approximately ten times the number of convolutional filters of the CNNs. Also, when comparing the ResNet-50 and ResNet-18 (see Fig. A.1 for ResNet-50), the larger network shows higher pruning potential.
4.2 Finding the Optimal Attributions for CNNs
Figure 4: Pruning CNNs models pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class. Results show a better sparsification-performance tradeoff for our optimized LRP composite compared to a heuristic (faithful) LRP composite, Yeom et al. [49] (details for each in Tab. A.2, Sec. 0.A.5, and Sec. 0.A.3.3) and random pruning.
Table 1: Results for pruning CNN models pre-trained on ImageNet. Top-PR demonstrates the highest pruning rate while keeping 95% of baseline accuracy. F, IG, and R in order indicate a faithful LRP composite, Integrated Gradients, and a random pruning baseline. Our results remark high scores for the area under the induced curve ($A^{\text{PR}}$) in pruning, overall improving the performance-sparsity tradeoff.
[…]
Models Ours F [49] IG R Ours F [49] IG R
VGG-16 0.61 0.58 0.58 0.43 0.4 26 23 21 7 6
VGG-16-BN 0.61 0.57 0.53 0.50 0.35 28 23 17 16 5
ResNet-18 0.69 0.68 0.57 0.53 0.37 41 40 18 20 5
ResNet-50 0.72 0.71 0.67 0.47 0.37 45 45 27 15 5
Motivated by the previous Sec. 4.1, we in the following simulate a setting with over-parameterized networks that allows us to measure more significant differences between methods. Specifically, we restrict the data domain to three ImageNet classes and evaluate pruning using 20 different random seeds. Experiments using toy models in the work of [49] showed that ten (or more) reference samples are already well suited for estimating the overall relevances of network components (as used in Eq. 1). We validated their finding also for the ResNet-18 model, as depicted in Fig. A.2.
Whereas we begin with optimizing LRP for CNN pruning, we follow up with transformers in the next section. Regarding CNNs, we can strongly reduce the accuracy-sparsity tradeoff when pruning the convolution filters with our method based on optimizing LRP attributions compared to other baselines, e.g., the heuristically chosen variant of LRP used by [49] (see Sec. 0.A.3.3), as also illustrated in Fig. 4 and Tab. 1. Extensive hyperparameter search reveals that across different CNN architectures, simple hyperparameter settings exist that are well-suited for pruning in general. One such candidate is the LRP-$\epsilon$ rule (see Eq. A.3) applied for all convolutional layers. We refer to Tab. A.2 for a list of the best performing hyperparameter settings. Interestingly for CNNs, LRP composites known to lead to faithful explanations (in the suggested context of [38, 5]) optimized by [25], are also effective in pruning. One such composite (detailed in Sec. 0.A.5) is also shown in Fig. 4 and denoted as “Faithful LRP”. Moreover, our results indicate a larger amount of unused structures across architectures with shortcut connections, i.e., the ResNet models. As intermediate features can be passed through the short-cuts, layers have the potential to be not used at all.
It is further not surprising that faithful LRP composites w.r.t. input explanations also perform well in pruning latent structures. Commonly, faithfulness is measured by “deleting” or perturbing input features and measuring the effect on the model’s prediction [38, 22]. A highly (or lowly) relevant feature is assumed to result in a high (or low) change in model output. As such, attribution-based pruning in combination with measuring the model performance resembles a faithfulness evaluation scheme in latent space. The more faithful the attributions, the smaller should be the degradation of model performance. However, later, in Sec. 4.3, we can also observe ineffective pruning with an attributor that is known to be faithful in input space.
It is to note that the LRP-$\varepsilon$ rule is known to result in noisy (and not necessarily faithful [38]) input attributions for DNNs, but performs very well for attributing latent components, as visible in Tab. A.2. We hypothesize that this results from fewer noise in intermediate layers [6] and the fact that we aggregate attributions for each component (e.g., channel) which further reduces noise, as also proposed for gradients for the GradCAM method [39]. High latent faithfulness with noisy input attributors (LRP-$\varepsilon$ or gradient times input [40]) has also been observed in [13, 14].
4.3 Finding the Optimal Attributions for Vision Transformer
Figure 5: Attribution-based pruning using a different number of reference samples (per class) to estimate the importance of attention heads (left) or neurons in linear layers (right) of the ViT-B-16. This experiment has been conducted for 20 different random seeds. For the propagation of LRP, LRP-$\epsilon$ has been set as our parameter for all layers (Sec. 3.3), and w.r.t. the attribution of softmax operations (Sec. 0.A.4).
Figure 6: Pruning a ViT model pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class. Results show a better sparsification-performance tradeoff for our optimized LRP composite compared to a heuristic (faithful) LRP composite, Yeom et al. [49] (details for each in Tab. A.2, Sec. 0.A.5, and Sec. 0.A.3.3, respectively), LRP-$\epsilon$, and random pruning.
Table 2: Results for pruning the ViT-B-16 pre-trained on ImageNet. Top-PR demonstrates the highest pruning rate while keeping 95% of baseline accuracy. F, IG, and R in order indicate a faithful LRP composite, Integrated Gradients, and a random pruning baseline. Our results remark high scores for the area under the induced curve ($A^{\text{PR}}$) in pruning, overall improving the performance-sparsity tradeoff.
[…]
Models Ours F LRP-$\epsilon$ [49] IG R Ours F LRP-$\epsilon$ [49] IG R
Linear Layers 0.87 0.75 0.83 0.77 0.70 0.59 57 51 49 48 33 26
Attention Heads 0.85 0.78 0.85 0.74 0.75 0.76 66 51 64 44 39 47
There are two structures of interest in the Vision Transformers for pruning in literature: attention heads and linear layers [45, 27]. As observed in Sec. 4.1, the ViT model shows a higher degree of over-parameterization, possibly due to the abundance of neurons in the linear layers, ultimately highlighting the possible value for pruning. We now re-investigate the number of necessary reference samples to robustly estimate the relevance of a model component of a ViT-B-16 model. Interestingly, unlike CNNs, according to Fig. 5, there is no deviation in pruning reliability and stability by using different numbers of reference samples. Consequently, our experiments have been applied with exact same settings as in Sec. 4.2.
Our results from Fig. 6 again confirm an improvement of our optimized attribution-based pruning scheme (best LRP composites shown in Tab. A.1) compared to the previous work of [49]. Adopting the generally reliable composite of LRP-$\epsilon$ obtained from CNNs (as in Fig. 4 and Tab. A.2) is again a promising option. Nonetheless, unlike for CNNs, a recently proposed faithful LRP composite [3] designed for the ViT-B-16 model (detailed in Sec. 0.A.5), is not ideal for pruning. This finding, and our experiment in Sec. 0.F.2 show that optimizing an attributor for two different contexts of faithfulness (input or latent space/pruning) (Sec. 3.3) does not necessarily lead to an attributor that attributes faithfully in both input and latent space. The poorer performance of heuristically tuned LRP from [49] (Sec. 0.A.3.3) compared to random when pruning attention heads, underscores both the greater challenge of pruning this structure and the significance of our proposed optimization procedure.
Figure 7: Pruning a ViT with the aim to retain a high accuracy for detecting ImageNet corgi classes (“Pembroke Welsh” and “Cardigan Welsh”). We show, how explanation heatmaps for a corgi prediction change when increasing the pruning rate (bottom). Initially, most positive relevance lies on the corgi’s head, with few negative relevance on a cat sitting next to the dog. As expected, when increasing the pruning rate using our optimized LRP composite, the background features disappear before corgi-related features are perturbed. When naively using the composite proposed in [49], or performing random pruning, heatmaps change much earlier (and more randomly), indicating model degradation. In fact, we can measure a high correlation of 0.99 between heatmap change and confidence loss (top right). Surprisingly, the unoptimized LRP composite performs even worse than random pruning, stressing the need to optimize attribution methods before applying them heuristically.
4.4 How Pruning Affects Model Explanations
In addition to keeping track of the model performance, the field of XAI encourages us in the following to investigate the model behavior by observing explanation heatmaps. Concretely, we expect in the ideal pruning scheme, that heatmaps change as late as possible when increasing the pruning rate. This reflects that the task-relevant components are retained as long as possible.
As an illustrative example, we prune the attention heads of the ViT-B-16 model pre-trained on ImageNet with the aim to predict ImageNet corgi classes (“Pembroke Welsh” and “Cardigan Welsh”) as shown in Fig. 7. As a quantitative measure, we compute the cosine similarity between the original heatmap (using the recently proposed composite of [1]) and the heatmap of the pruned model for different pruning rates over the validation set.
On the one hand, we can see that random pruning or unoptimized attribution-based pruning leads to a much earlier change in heatmaps compared to our optimized LRP-based approach. The heatmaps indicate that irrelevant features are removed first as the pruning rate is increased (e.g., “cat” features in Fig. 7 get disregarded before corgi-related features in the process). However, a random pruning approach might wrongly omit an important structure first, as heatmap changes of “mouth” and “ear” features can be seen in Fig. 7. On the other hand, as can be expected, the change in heatmap similarity highly correlates with the model confidence, resulting in a Pearson correlation coefficient of 0.99.
5 Conclusion
In this work, we propose a general framework for post-hoc DNN pruning that is based on using and optimizing attribution-based methods from the field of eXplainable Artificial Intelligence for more effective pruning. For our framework, the method of LRP is well-suited by offering several hyperparameters to tune attributions. When applying our framework, we can strongly reduce the performance-sparsity tradeoff of CNNs and especially ViTs compared to previously established approaches. Vision transformers are on the one hand more sensitive towards hyperparameters, and also show higher over-parameterization. Overall, using local XAI methods for pruning irrelevant model components demonstrates high potential in our experiments.
Appendix
Appendix 0.A Attribution Methods
As it has been discussed in Sec. 3.1, backpropagation methods are suitable for our proposed framework. We focus mostly on LRP and it will be compared with Integrated Gradient as a commonly used gradient-based explainer.
0.A.1 Integrated Gradient
As an improvement to pure gradient, Integrated Gradient [44] interpolates input $x$ for $m$ steps and computes their gradients sequentially.
[…]
$\approx(x-x^{\prime})\sum_{k=1}^{m}\frac{\partial f_{j}(x^{\prime}+\frac{k}{m}\times(x-x^{\prime}))}{\partial x}\times\frac{1}{m}$ (A.2)
Unlike LRP, there is no parameter to be tuned in this method. $m$ from equation Eq. A.1 only serves as an approximation factor, with higher steps indicating more precise computation of the integral. For the experiments conducted in this paper, we interpolate each input for $20$ steps.
0.A.2 Variants of LRP
In addition to the vanilla rule of LRP demonstrated in 6, [5, 33] later on proposed different rules as an extension of vanilla rule designed to serve different purposes. Most common extension of LRP will be discussed based on the layer type they target.
0.A.3 Tackling Linear and Convolution Layers
As a side note, in the propagation process of LRP, linear layers of DNNs such as fully connected, and convolution layers, are treated similarly.
0.A.3.1 LRP-$\epsilon$
The most fundamental problem of rule 6 causing computational instability, is division by zero which takes place when a neuron is not activated at all ($z_{j}=0$). LRP-$\epsilon$ with $\epsilon\in\mathbb{R}$ as a stabilizer parameter, was proposed to tackle this problem so that the denominator never reaches zero:
$\centering R_{i\leftarrow j}=\frac{z_{ij}}{z_{j}+\epsilon\text{sign}(z_{j})}R_{j}\@add@centering$ (A.3)
As a side note, sign(0) = 1. Although the value of $\epsilon$ can be tuned, we set it to $61e-$ 6 in every use-case.
0.A.3.2 LRP-$\alpha\beta$
As described in figure Fig. 2 and Sec. 3.3, relevances either have a positive or negative sign, indicating whether they have contributed positively or negatively in favor of the decision-making. Depending on the use-case, we might be interested in having a higher emphasis on positive relevances rather than the negative ones or vice versa. LRP-$\alpha\beta$ rule with $\alpha+\beta=1$, lets us apply modifications in that regard.
$\centering R_{i\leftarrow j}=\bigr{(}\alpha\frac{z^{+}_{ij}}{z^{+}_{j}}+\beta\frac{z^{-}_{ij}}{z^{-}_{j}}\bigr{)}R_{j}\@add@centering$ (A.4)
0.A.3.3 LRP-$z^{+}$
An extreme case for the above formulation when it is required to ignore all negative relevance, is the combination of $\alpha=1$ and $\beta=0$, which is referred to as LRP-$z^{+}$ used by [49].
0.A.3.4 LRP-$\gamma$
Another derivation, with the same purpose of LRP-$\alpha\beta$ to control over the distribution of positive and negative relevance separately, has this form:
$\centering R_{i\leftarrow j}=\frac{z_{ij}+\gamma z_{ij}^{+}}{z_{j}+\gamma\sum_{k}z_{kj}^{+}}R_{j}\@add@centering$ (A.5)
Evidently, in case of $\gamma=\infty$, LRP-$\gamma$ will be equivalent to LRP-$z^{+}$. The value of $\gamma$ can also be tuned in the optimization process. However, to prevent further computational costs unlike results from [2] in Sec. 0.A.5, we refrain from tuning it.
0.A.4 Tackling Attention and Softmax Non-linearity
The key component of the attention module [45], consists of two equations:
$\textbf{A}=\text{softmax}\left(\frac{\textbf{Q}\cdot\textbf{K}^{T}}{\sqrt{d_{k}}}\right)$ (A.6)
[…]
where ($\cdot$) denotes matrix multiplication, $\textbf{K}\in\mathbb{R}^{b\times s_{k}\times d_{k}}$ is the key matrix, $\textbf{Q}\in\mathbb{R}^{b\times s_{q}\times d_{k}}$ is the queries matrix, $\textbf{V}\in\mathbb{R}^{b\times s_{k}\times d_{v}}$ the values matrix, and $\textbf{O}\in\mathbb{R}^{b\times s_{k}\times d_{v}}$ is the final output of the attention mechanism. $b$ is the batch dimension including the number of heads, and $d_{k},d_{v}$ indicate the embedding dimensions, and $s_{q},s_{k}$ are the number of query and key/value tokens.
0.A.4.1 CP-LRP
In the extension of LRP from [4], it was proposed to regard the attention matrix (softmax output) A in Eq. A.7 as constant, attributing relevance solely through the value path by stopping the relevance flow through the softmax. Consequently, the matrix multiplication Eq. A.7 can be treated with the $\varepsilon$-rule Eq. A.3.
0.A.4.2 AttnLRP
The recent work of [2] extends LRP to Large Language Models and Vision Transformers by proposing to linearize the softmax function in Eq. A.6 with a Deep Taylor Decomposition [34] which leads to substantially improved attributions in the natural language domain. For that, the softmax non-linearity Eq. A.6 is attributed as follows:
$R^{l}_{i}=x_{i}(R^{l+1}_{i}-s_{i}\sum_{j}R^{l+1}_{j})$ (A.8)
where $s_{i}$ denotes the $i$-th output of the softmax. In addition, the matrix multiplication Eq. A.7 is attributed with
$R_{ji}^{l-1}(\textbf{A}_{ji})=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R^{l}_{jp}}{2\ \textbf{O}_{jp}+\epsilon}$ (A.9)
Finally, in order to mitigate noise in the attributions for Vision Transformers, the authors [2] propose to apply the LRP-$z^{+}$ on the Deep Taylor linearization of the softmax non-linearity Eq. A.6:
$R_{i}^{l-1}=\sum_{j}(\textbf{J}_{ji}\ x_{i})^{+}\frac{R_{j}^{l}}{\sum_{k}(\textbf{J}_{jk}\ x_{k})^{+}+\tilde{b}_{j}^{+}}$ (A.10)
where $J$ is the Jacobian of the softmax computed at the current point $x$ and $\tilde{b}$ is the bias term of the linearization. For more details, please refer to [2].
0.A.5 Faithful LRP Composites
The works of [25, 2] introduced 2 different LRP composites as faithful explainer [38]. For CNNs, [25] suggested applying LRP-$z^{+}$ to FCLs and LRP-$\epsilon$ to other remaining layers (including HLLs, MLLs, LLLs).
On the other hand, experiments from [2] over the layers of transformers (differentiated based on layer types) demonstrated simple LRP-$\epsilon$ (Eq. A.3) generates faithful explanation for Large Language Models (LLMs). However, unlike LLMs, the composite of Tab. A.1 seems to be more suitable for ViTs.
Table A.1: Faithful LRP composite for Vision Transformers. LinearInputProjection and LinearOutputProjection are the linear layers generating $Q$, $K$, $V$, and $O$ inside attention-module of Eq. A.7 and Eq. A.6. The values of $\gamma$ for applying LRP in Convolution and Linear layers are $0.25$ and, $0.05$ respectively.
Conv. Linear LinearInputProjection LinearOutputProjection Softmax
$\gamma$ $\gamma$ $\epsilon$ $\epsilon$ Attn-LRP$(z^{+})$
Appendix 0.B Spatial Dimension of To-Be-Pruned Components
In Eq. 1, $R_{p}(x_{i})$ presented the relevance of component $p$ for reference sample $x_{i}$. However, it is crucial to discuss the spatial dimensionality of each component depending on its type because the relevance of a component gets aggregated over its dimensions. Convolution filters of CNNs have the size of the tuple $(h,w,d)$ indicating the height, weight, and depth of a filter. Thus, for a batch of size $n$ and convolution filter $p$, Tab. A.1 will be transformed to:
$\centering\bar{R_{p}}=\sum_{i}^{n}\sum_{j}^{h}\sum_{l}^{w}\sum_{r}^{d}R_{(p,j,l,r)}({x_{i}})\@add@centering$ (A.11)
Unlike linear layers of CNNs which simply follow Eq. 1, an extra token axes $(t)$ for neurons inside linear layers of transformers resulting in this modification:
$\centering\bar{R_{p}}=\sum_{i}^{n}\sum_{j}^{t}R_{(p,j)}({x_{i}})\@add@centering$ (A.12)
Attention heads of Eq. A.6, have two extra axes $(d_{q},d_{k})$ regarding as Query, and Key. Aggregation of relevance follows below formula:
$\centering\bar{R_{p}}=\sum_{i}^{n}\sum_{j}^{d_{q}}\sum_{l}^{d_{k}}R_{(p,j,l)}({x_{i}})\@add@centering$ (A.13)
However, in cases of computing the magnitude of relevance, aggregation should follow this structure:
$\centering\bar{R_{p}}=\sum_{i}^{n}\sum_{j}^{d_{q}}\big{|}\sum_{l}^{d_{k}}R_{(p,j,l)}({x_{i}})\big{|}\@add@centering$ (A.14)
Appendix 0.C Additionally on Overparameterization
Similar to Fig. 3, Fig. A.1 demonstrates the effect of overparameterization additionally on VGG-16 and ResNet-50 signifying again over-use of parameters in simpler tasks. Similarly in this case, architectures with skip connections show more potential to be overparameterized.
Figure A.1: With settings similar to Fig. 3, few-shot pruning demonstrates overparameterization exceeds, especially in cases of simpler tasks. Additionally, Relevance denoted in the legend is computed by LRP framework with $\epsilon$ rule (see Sec. 0.A.2) applied across all layers.
Appendix 0.D On Optimization: From Bayesian to Grid Search
To find an optimal LRP composite for pruning, our approach takes place by discovering prospective parameters (Sec. 3.3) via Bayesian Optimization [16] (with Gaussian Process regressor [47] as a surrogate model back-bone), followed by Grid Search on the reduced parameter space. This seems to be a more effective solution rather than naively applying Grid Search over the whole parameters when we can reduce from $50\%$ up to $90\%$ of our search space approximately.
Appendix 0.E On Number of Reference Samples
Fig. A.2 indicates and validates the findings of work [49] that using minimally 10 reference samples results in stable pruning for CNNs, as no more impact can be observed when using more samples. The fewer used samples per class lead to a lower pruning rate.
Figure A.2: Tracking changes of pruning rate for a ResNet-18 pre-trained on ImageNet based on the variable number of reference samples to conduct pruning via LRP for 20 different random seeds. We used LRP-$\epsilon$ over all layers.
Appendix 0.F Optimized Composites
0.F.1 Top Composites for Pruning
In later tables, based on the conducted optimization, we demonstrate top composites for pruning on CNNs (Tab. A.2), ViT-B-16 (Tab. A.3) and propose a general composite with persuasive performance to use over each architecture type individually. Later, in Tab. A.4 we discuss which parameters are more important to be tuned.
[…]
VGG-16 $\epsilon$ $\epsilon$ $\epsilon$ $\epsilon$ True 0.608 26
VGG-16-BN $\epsilon$ $\epsilon$ $\epsilon$ $\epsilon$ True 0.609 28
ResNet-18 $\epsilon$ $\epsilon$ $\epsilon$ $\alpha^{2}\beta^{1}$ True 0.691 41
ResNet-50 $\epsilon$ $\epsilon$ $\epsilon$ $\epsilon$ True 0.718 45
Across All $\epsilon$ $\epsilon$ $\epsilon$ $\epsilon$ True 0.656
[…]
Attention Heads $\epsilon$ $\epsilon$ $z^{+}$ $\alpha^{2}\beta^{1}$ $\epsilon$ True 0.851 57
Linear Layers $\epsilon$ $\alpha^{2}\beta^{1}$ $\gamma$ $\gamma$ CP-LRP True 0.871 66
Across All $\epsilon$ $\epsilon$ $\epsilon$ $z^{+}$ $z^{+}$ True 0.851
Table A.4: The importance of tuning an individual parameter that is involved in constructing LRP composite and pruning framework, can be assessed by computing the deviation in performance when that specific parameter is altered while others remain constant. The most influential ones per architecture are displayed in bold font.
Model LLL MLL HLL FC Softmax MAG
VGG-16 0.028 0.075 0.047 0.030 0.120
VGG-16-BN 0.009 0.062 0.072 0.025 -
ResNet-18 0.019 - 0.39 0.17 -
ResNet-50 0.059 0.72 0.070 0.017 0.205
ViT-B-16 (Linear Layers) 0.102 0.077 0.05 0.012 0.006 -
ViT-B-16 (Attention Heads) 0.012 0.048 0.025 0.029 0.012 0.061
0.F.2 Evaluating the Composite
As described in Sec. 3.3, we proposed our pruning framework as a means of evaluating attribution methods in terms of aligning with the actual importance of latent components. However, it is noted that composites highly scored by this criterion, are not necessarily performant in generating explanation heatmaps. Our results in Sec. 4.3 exhibited that LRP composite (Sec. 0.A.5) suggested by [2] is not the most suitable option for pruning. However, qualitative and quantitative experiment (based on [6, 5, 2]) in Fig. A.3 demonstrates that our best composite (see Tab. A.1) is not proper to render explanation heatmap. All in all, LRP explainers tuned to perform well in pruning, do not necessarily induce reliable heatmaps.
Figure A.3: Qualitative experiment of ViT-B-16 on the left shows that the composite from [2] generates more localized and less noisy explanations in the input-space rather than ours, proposed in Tab. A.1. A quantitative experiment of evaluating attribution methods on right side of the figure validates this finding. The higher area under the curve induces a more faithful heatmap. This experiment took place by using around $1000$ samples from ImageNet and $100$ perturbation steps (see work of [8] for the description of this metric).
Appendix 0.G Layer-wise Relevance Flow
Later figures (A.4, A.5, A.6, A.7, A.8, A.9) partially elucidate the pruning strategy by illustrating relevance magnitude of components in a layer-wise fashion.
Figure A.4: There are two linear layers inside 12 Encoder layers of ViT-B-16. Magnitude of LRP-$\epsilon$ relevance shows higher magnitudes as we approach lower layers. Interestingly, the second layers in each Encoder block has generally higher relevance values, which can be expected, as here 512 instead of 2048 neurons (as in the previous layer) are included.
Figure A.5: 144 attention heads are distributed in 12 Encoder layers of ViT-B-16 highlight higher relevance in mid-level layers, corresponding to higher importance in decision-making.
Figure A.6: Generally higher relevance for mid to low level convolution filters. There are in total, 4800 convolution filters in ResNet-18 distributed in 20 layers.
Figure A.7: There are in total, 26560 convolution filters in ResNet-50 distributed in 53 layers. Higher relevance magnitude in first layers generally stems from having a few number of convolution filters there.
Figure A.8: There are in total, 4224 convolution filters in VGG-16 distributed in 13 layers. Higher relevance magnitude is visible at mid to lower-level convolution filters.
Figure A.9: Similar to Fig. A.8 there are in total, 4224 convolution filters in VGG-16-BN distributed in 13 layers. Higher relevance magnitude is visible at mid to lower-level convolution filters.
Appendix 0.H Change in Behaviour
In later figures (A.10, A.11, A.12), we qualitatively demonstrate examples of pruning based on 3 randomly chosen classes and trace changes in the explanation heatmap of the pruned model.
Figure A.10: Changes in ResNet-18 decision-making for domain restriction of 3 randomly chosen classes (Box Turtle, Alligator Lizard, and Moving Van) via LRP heatmaps (using composite described in Sec. 0.A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and background despite considering the main object itself. Heatmaps of the model with high pruning rates demonstrate an effect caused by a subsampling shortcut connection becoming apparent.
Figure A.11: Changes in ResNet-50 decision-making for domain restriction of 3 randomly chosen classes (Flamingo, Gordon Setter, and Snowplow) via LRP heatmaps (using composite described in Sec. 0.A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and unlike ResNet-18 ignores the main object. The shortcut connections of ResNet architecture result in checkerboard pattern exhibited in the heatmaps.
Figure A.12: Changes in VGG-16-BN decision-making for domain restriction of 3 randomly chosen classes (Digital Clock, Patio, Terrace, and Prayer Rug) via LRP heatmaps (using composite described in Sec. 0.A.5) indicate lower pruning capabilities of this architecture compared to ResNet-18 and ResNet-50.

### v2

@@
-
+Pruning By Explaining Revisited: Optimizing Attribution Methods to Prune CNNs and Transformers
 Abstract
@@
 We extend the current state by proposing to explicitly optimize hyperparameters of attribution methods for the task of pruning, and further include transformer-based networks in our analysis.
-Our approach yields higher model compression rates of large transformer- and convolutional architectures (VGG, ResNet, ViT) compared to previous works, while still attaining high performance on ImageNet classification tasks.
+Our approach yields higher model compression rates of large transformer and convolutional architectures (VGG, ResNet, ViT) compared to previous works, while still attaining high performance on ImageNet classification tasks.
 Here, our experiments indicate that transformers have a higher degree of over-parameterization compared to convolutional neural networks.
 Code is available at https://github.com/erfanhatefi/Pruning-by-eXplaining-in-PyTorch.
-Keywords:
-Explainable AI Pruning Attribution Optimization
+Keywords Explainable AI  $\cdot$ Pruning  $\cdot$ Attribution Optimization
 1 Introduction
-Figure 1: We propose a pruning framework based on optimizing attribution methods from the field of XAI.
-Compared to random pruning, pruning the least relevant structures first (“relevant” according to an XAI attribution method of choice, and indicated by red color), has been shown to result in an improved performance-sparsity tradeoff (simplified illustration depicted).
-By optimizing attribution methods specifically for pruning, we can reduce the tradeoff even further.
-In recent years, Deep Neural Networks (DNNs) have been growing increasingly larger, demanding more computational resources and memory.
+In recent years, Deep Neural Networks (DNNs) have been growing larger increasingly, demanding ever more computational resources and memory.
 To address these challenges, several efficient architectures, such as MobileNet [22] or EfficientFormer [28], have been proposed to reduce computational costs.
-However, the gain in efficiency comes at the cost of performance, as inherently efficient architectures struggle to keep pace with the recent surge in high-performing transformer models.
+However, the gain in efficiency comes at the cost of performance, as inherently efficient architectures struggle to keep pace with the recent surge in high-performing but costly transformer models.
 One widely adopted approach within the community of efficient Deep Learning is quantization [17, 49], which involves compressing model parameters or features by reducing the number of bits used for representation, potentially followed by re-training to regain lost model performance.
@@
 Earlier works have shown that structures may be chosen according to, e.g., parameter magnitudes [18, 27], functional redundancy [16], sensitivity [47, 31], or importance in the decision-making process [48, 6].
-Especially the latter approach has been promising, utilizing tools from the seemingly unrelated field of XAI that originally seeks to explain the model reasoning process to human stakeholders.
+Especially the latter approach has been promising, utilizing tools from the seemingly unrelated field of eXplainable Artificial Intelligence (XAI) that originally seeks to explain the model reasoning process to human stakeholders.
+Figure 1: We propose a pruning framework based on optimizing attribution methods from the field of XAI.
+Compared to random pruning, pruning the least relevant structures first (“relevant” according to an XAI attribution method of choice, and indicated by red color), has been shown to result in an improved performance-sparsity tradeoff (simplified illustration depicted).
+By optimizing attribution methods specifically for pruning, we can reduce the tradeoff even further.
 Backpropagation-based attribution methods, particularly Layer-wise Relevance Propagation (LRP) [4, 32, 33], play a crucial role in this context.
@@
 These scores provide a quantitative basis for identifying (ir-)relevant structures and subgraphs within the model that are (un-)essential for making accurate predictions, thus guiding the pruning process.
-While prior attribution-based pruning [49] achieved remarkable results, it is restricted to heuristically chosen LRP hyperparameters and limited to specific Convolutional Neural Networks (CNNs) and toy models.
+While prior attribution-based pruning [48] achieved remarkable results, it is restricted to heuristically chosen LRP hyperparameters and specific Convolutional Neural Networks (CNNs).
 Following up upon the work [34] that optimized LRP w.r.t.
-specific explainability metrics, we argue that LRP or any other attribution method can be optimized specifically for the goal of more effective pruning (see Fig.
+specific explainability metrics, we argue that LRP or any other tunable attribution method can be optimized specifically for the goal of more effective pruning (see Fig.
 1).
@@
   4. 4.
-    revealing that attributions optimized for explanations are not necessarily the best for pruning. Notably, simple hyperparameter settings show high performance across all tested architectures.
+    revealing that attributions optimized for explanations are not necessarily the best for pruning.
 2 Related Work
 In the following, we introduce related works in the field of XAI, efficient Deep Learning and the intersection between both.
-Explainability and Local Feature Attribution Methods
-Research in local explainability led to a plethora of methods [30, 37], commonly resulting in local feature attributions quantifying the importance of input features in the decision-making process.
+Explainability and Local Feature Attribution Methods.
+Research in local explainability led to a plethora of methods (e.g., [29, 36]), commonly resulting in local feature attributions quantifying the importance of input features in the decision-making process.
 These attributions are often shown in the form of heatmaps in the vision domain.
@@
 We follow up on this observation, and specifically optimize XAI w.r.t.
-pruning.
-Pruning of Deep Neural Networks
+the task of pruning.
+That is, we add Neural Network (NN) pruning as an XAI evaluation criterion to optimize for.
+Pruning of Deep Neural Networks.
 For pruning CNNs, either individual (kernel) weights or whole structures, e.g., filters of convolution layers or neurons can be recognized as candidates for pruning [20].
@@
 In order to prune such structures, several criteria have been proposed to indicate which components are best suited to be removed, retaining performance as best as possible.
-The work of [19] suggests pruning parameters based on weight magnitudes, offering a computationally-free criterion.
+The work of [18] suggests pruning parameters based on weight magnitudes, offering a computationally free criterion.
 Alternatively, in [10], the authors propose to prune neurons based on their activation patterns.
-However, recent work [7] has shown that weight or activation magnitudes do not necessarily reflect a component’s contribution during the inference process, e.g., also a small weight has the potential to be very relevant for the predication of a class.
-Explainability For Efficient Deep Learning
+However, recent work [6] has shown that weight or activation magnitudes do not necessarily reflect a component’s contribution during the inference process, e.g., also a small weight has the potential to be very relevant for the prediction of a class.
+Explainability For Efficient Deep Learning.
 The work of [48] introduces a novel pruning criterion based on XAI, and proposes to use latent relevance values extracted from the attribution method of LRP [4, 33] to assign importance scores to network structures.
-By taking into account how structures are used during the inference process, [49] can effectively improve the efficiency-performance tradeoff.
+By taking into account how structures are used during the inference process, the work of [48] can effectively improve the efficiency-performance tradeoff.
 Later, the works of [6, 42] also illustrate the value of XAI methods for network quantization.
@@
 Notably, LRP is model-specific and thus restricted to compatible architectures.
-In this work, we also include recent LRP extensions to transformer architectures, and observe, that simple LRP settings commonly result in high pruning performances.
+In this work, we also include recent LRP extensions to transformer architectures, and observe, that a specific LRP rule commonly result in high pruning performances.
 3 Methods
@@
 For our structured pruning framework, we view a DNN as a collection of $p$ (interlinked) components $\Psi=\{\psi_{1},\dots,\psi_{p}\}$, that can correspond to, e.g., whole layers, (groups of) neurons, convolutional filters or attention heads.
-We further assume access to an attribution method that generates attribution scores (relevance) $R_{\psi_{k}}({x_{i}})$ of component $\psi_{k}\in\Psi$ for the prediction of a sample $x_{i}$.
+We further assume access to an attribution method that generates attribution scores (relevance scores) $R_{\psi_{k}}({x_{i}})$ of component $\psi_{k}\in\Psi$ for the prediction of a sample $x_{i}$.
 The overall relevance of $\psi_{k}$ is then estimated through the mean relevance over a set of reference samples $\mathcal{X}_{\text{ref}}=\{x_{1},x_{2},\dots,x_{n_{\text{{ref}}}}\}$ as
-Figure 2: Attribution-based pruning workflow: Firstly, the relevant model structures are found by explaining a set of reference samples.
+Figure 2: Attribution-based pruning workflow: Firstly, the relevant model structures are identified by explaining a set of reference samples.
 The attribution method of choice (here LRP) highlights the components and paths in the network which positively and negatively contribute to the decision-making.
-Positive and negative relevances are indicated by red and blue color respectively, and components with white color indicate no relevance.
+Positive and negative relevances are indicated by red and blue color respectively, and components with white color indicate zero or low relevance.
 Removing structures that receive the least relevance results in a sparser subnetwork, which performs significantly better than after random pruning.
@@
 Credit: Nataba/iStock.
-$\bar{R}_{\psi_{k}}=\frac{1}{n_{\text{ref}}}\sum_{i=1}^{n_{\text{ref}}}R_{\psi_{k}}({x_{i}})\,.$ (1)
-Ultimately, we collect relevances for all components via the set $\mathcal{R}$, given as
-=$\mathcal{R}=\{\bar{R}_{\psi_{1}},\bar{R}_{\psi_{2}},\dots,\bar{R}_{\psi_{p}}\}\,,$ (2)
+$\bar{R}_{\psi_{k}}=\frac{1}{n_{\text{ref}}}\sum_{i=1}^{n_{\text{ref}}}R_{\psi_{k}}({x_{i}})$ (1)
+We collect relevance scores for all components via the set $\mathcal{R}$, given as
+=$\mathcal{R}=\{\bar{R}_{\psi_{1}},\bar{R}_{\psi_{2}},\dots,\bar{R}_{\psi_{p}}\}$ (2)
 which, in turn, allows to define a pruning order for the model components. Specifically, the indices $c$ for the components to be pruned up to the $q$-th place are given by
@@
 2.
-Notably, previous work [49] demonstrates that LRP is an effective method for attributing latent structures, as used for Eq.
-1.
+Notably, previous work [48] demonstrates that LRP is an effective method for attributing latent structures.
 LRP further offers several hyperparameters to tune, which makes it a versatile choice for our framework.
@@
 $\centering f(x)=f^{L}\circ\dots\circ f^{l}\circ f^{l-1}\circ\dots\circ f^{1}(x)\@add@centering$ (5)
-Beginning with an initial relevance score $R_{j}^{L}$ at output $j$ of layer $f^{L}$, the score is layer-by-layer redistributed to its input variables depending on the contribution from the input variables to the output value.
+Beginning with an initial relevance score $R_{j}^{L}$ at output $j$ of layer $f^{L}$ (usually set as $f_{j}^{L}$ for an output of choice $j$), the score is layer-by-layer redistributed through all latent structures to its input variables depending on the contribution from the these units to the output value:
 Given a layer, we consider its pre-activations $z_{ij}$ mapping inputs $i$ to outputs $j$ and their aggregations $z_{j}=\sum_{i}z_{ij}$.
@@
 Then, LRP distributes relevance quantities $R_{j}^{l}$ received from upper layers towards lower layers proportionally to the relative contributions of $z_{ij}$ to $z_{j}$, i.e.,
-$\centering R_{i\leftarrow j}^{(l-1,l)}=\frac{z_{ij}}{z_{j}}R_{j}^{l}\,.\@add@centering$ (6)
+$\centering R_{i\leftarrow j}^{(l-1,l)}=\frac{z_{ij}}{z_{j}}R_{j}^{l}\@add@centering$ (6)
 In other words, the relevance message $R_{i\leftarrow j}^{(l-1,l)}$ quantifies the contribution of neuron $i$ at layer $l-1$, to the activation of neuron $j$ at layer $l$.
 To obtain the contribution of neuron $i$ to all upper layer neurons $j$, all incoming relevance messages $R_{i\leftarrow j}^{(l-1,l)}$ are losslessly aggregated as
-$R_{i}^{l-1}=\sum_{j}R_{i\leftarrow j}^{(l-1,l)}.$ (7)
-This process ensures the property of relevance conservation between adjacent layers:
+$R_{i}^{l-1}=\sum_{j}R_{i\leftarrow j}^{(l-1,l)}$ (7)
+This process ensures relevance conservation between adjacent layers:
 $\sum_{i}R^{l-1}_{i}=\sum_{i,j}R^{(l-1,l)}_{i\leftarrow j}=\sum_{j}R^{l}_{j}$ (8)
 which guarantees that the sum of all relevance in each layer stays the same.
-When a group of neurons performs the same task within a convolutional channel or attention head, it is beneficial to aggregate the total relevance of the entire group into a single relevance score.
+When a group of neurons performs the same task as within a convolutional channel or attention head, it is beneficial to aggregate the total relevance of the entire group into a single relevance score.
 This aggregation process helps in simplifying the analysis and interpretation of the model’s behavior by focusing on the collective relevance of each convolutional channel or attention head, rather than examining individual neurons.
-Further discussions of component-wise aggregation are given in Appendix 0.B.
+Further discussions of component-wise aggregation are given in Appendix B.
 3.3 Tuneable Hyperparameters of LRP
-Within the LRP framework, various rules (as detailed in Sec. 0.A.2) have been proposed to tune relevance computation for higher explainability.
-Further, several works have shown that using different rules for different architecture parts (referred to as a “composite”) [33, 25, 2] is favorable for explanations as it improves robustness and faithfulness.
+Within the LRP framework, various rules (as detailed in Sec. A.2) have been proposed to tune relevance computation for higher explainability.
+Further, several works have shown that applying different rules to variable parts of a network (referred to as a “composite”) [32, 24, 2] enhances faithfulness and robustness of explanations.
 We follow up on this idea, and perform a large-scale analysis over a variety of rules for different network parts.
-Inspired by the tuning settings of [35], we split a network into four parts:
+Inspired by the tuning configurations of [34], we split a network into four parts:
   1. 1.
@@
     and the remaining hidden layers between HLL and LLL, denoted as Mid-Level hidden Layers (MLL)
-For each group, a specific rule can be chosen, which together ultimately form the LRP composite.
-We refrain from treating each individual layer as a separate group due to the increased computational expenses associated with this setting.
+For each group, a specific rule set can be chosen, which together ultimately form the LRP composite.
+We refrain from treating each individual layer as a separate group due to the increased computational expenses associated with this setting during hyperparameter search.
 For transformer architectures, the works of [2, 3] proposed novel LRP rules to attribute softmax non-linearities in the attention modules (see Sec.
-0.A.2 for more details), encouraging us to also optimize the LRP rule for the softmax non-linearities, resulting in an additional hyperparameter here.
+A.2 for more details), encouraging us to also optimize LRP configurations for the softmax non-linearities, resulting in an additional hyperparameter choices here.
 It is important to note, that relevance values can be positive or negative.
-Components with positive relevance values indicate that they significantly contribute towards the class prediction.
+Components with positive relevance values indicate that they significantly contribute towards the explained class prediction.
 Conversely, components with negative relevance values diminish the classification score, effectively speaking against the predicted class, as visible in Fig.
@@
 1, i.e., $\bar{R}_{\psi_{k}}\mapsto|\bar{R}_{\psi_{k}}|$ for all $\psi_{k}\in\Psi$ if true.
-The work of [49] takes into account only positive attributions by their LRP rule choice (see LRP-$z^{+}$ in Sec.
-0.A.2).
-Our experiments in Appendix 0.F, however, reveal that the highest pruning rates are achieved by first pruning components with near zero relevance, as their minimal contribution ensures low impact on the overall model performance.
+The work of [48] takes into account and computes only positive attributions by their LRP rule choice (see LRP-$z^{+}$ in Sec.
+A.2).
+Our experiments in Appendix F, however, reveal that the highest pruning rates are achieved by computing both positive and negative relevance signals, and then first pruning components with near zero relevance, as their minimal contribution in any direction ensures low impact on the overall model performance.
 3.4 Hyperparameter Optimization Procedure
@@
 $\theta$.
-Hyperparameter search is performed via grid search and Bayesian optimization techniques, as detailed in Appendix 0.D.
+Hyperparameter search is performed via grid search and Bayesian optimization techniques, as detailed in Appendix D.
 4 Experiments
@@
 Notably, one of the effects that arise when solving for simple(r) tasks, is that we likely end up with an over-parameterized model as only a subset of very specialized latent features are necessary to solve the task, which makes pruning especially interesting in this case.
-Figure 3: Investigating over-parameterization in DNNs through attribution-based pruning (LRP-$\epsilon$ in blue, unoptimized in orange color) and random pruning (green color).
+Figure 3: Investigating over-parameterization in DNNs through attribution-based pruning (LRP-opt in blue, Integrated Gradient in orange color) and random pruning (green color).
 We compare pruning of all models w.r.t.
@@
 However, as the task becomes simpler (simulated by reducing the number of output classes to 100 or three in the evaluation), pruning rates can be increased to a much higher level without critical accuracy loss.
-For baseline comparisons, we utilize random pruning, Integrated Gradient [44] and our optimized LRP attributions (see Sec.
+For baseline comparisons, we utilize random pruning, Integrated Gradient [43] and our optimized LRP attribution (LRP-opt) (see Sec.
 4.2).
-We intentionally exclude naive weight magnitude or activation pruning since [49] has previously demonstrated that XAI-based methods outperform these traditional approaches.
+We intentionally exclude naive weight magnitude or activation pruning since [48] has previously demonstrated that XAI-based methods outperform these traditional approaches by a large margin.
 Notably, especially for simpler tasks, attribution-based pruning outperforms random pruning, as attributions are output-specific and allow identifying the sub-graph that is relevant for the restricted task, as also illustrated in Fig.
@@
 Compared to the previously discussed CNNs, the ViT-B-16 vision transformer seems to be more over-parameterized, as visible in Fig.
-3 (bottom).
+3 (bottom), and therefore easier to prune.
 Up to a pruning rate of 20 %, no meaningful accuracy loss is measured for detecting the 1000 ImageNet classes, irrespective of whether attention heads or neurons in linear layers are pruned.
-This might be due to the fact, that the vision transformer consists of about 46 thousand neurons, which is approximately ten times the number of convolutional filters of the CNNs.
+This might be due to the fact, that the vision transformer consists of about 46,000 neurons, which is approximately ten times the number of convolutional filters of the CNNs.
 Also, when comparing the ResNet-50 and ResNet-18 (see Fig.
@@
 A.2, Sec.
-0.A.5, and Sec.
-0.A.3.3) and random pruning.
+A.5, and Sec.
+A.3.3) and random pruning.
 Table 1: Results for pruning CNN models pre-trained on ImageNet.
@@
 Regarding CNNs, we can strongly reduce the accuracy-sparsity tradeoff when pruning the convolution filters with our method based on optimizing LRP attributions compared to other baselines, e.g., the heuristically chosen variant of LRP used by [48] (see Sec.
-0.A.3.3), as also illustrated in Fig.
+A.3.3), as also illustrated in Fig.
 4 and Tab.
@@
 One such composite (detailed in Sec.
-0.A.5) is also shown in Fig.
+A.5) is also shown in Fig.
 4 and denoted as “Faithful LRP”.
 Moreover, our results indicate a larger amount of unused structures across architectures with shortcut connections, i.e., the ResNet models.
-As intermediate features can be passed through the short-cuts, layers have the potential to be not used at all.
+As intermediate features can be passed through the short-cuts, while layers have the potential to be not used at all.
 It is further not surprising that faithful LRP composites w.r.t.
@@
 the attribution of softmax operations (Sec.
-0.A.4).
+A.4.2).
 Figure 6: Pruning a ViT model pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class.
@@
 A.2, Sec.
-0.A.5, and Sec.
-0.A.3.3, respectively), LRP-$\epsilon$, and random pruning.
+A.5, and Sec.
+A.3.3, respectively), LRP-$\epsilon$, Integrated Gradient, and random pruning.
 Table 2: Results for pruning the ViT-B-16 pre-trained on ImageNet.
@@
 Attention Heads 0.85 0.78 0.85 0.74 0.75 0.76 66 51 64 44 39 47
-There are two structures of interest in the Vision Transformers for pruning in literature: attention heads and linear layers [45, 27].
+Attention heads and linear layers are two structures of interest in the pruning literature for Vision Transformers [44, 26].
 As observed in Sec.
 4.1, the ViT model shows a higher degree of over-parameterization, possibly due to the abundance of neurons in the linear layers, ultimately highlighting the possible value for pruning.
-We now re-investigate the number of necessary reference samples to robustly estimate the relevance of a model component of a ViT-B-16 model.
+We now re-investigate the number of reference samples required to robustly estimate the relevance of a component in the ViT-B-16 model.
 Interestingly, unlike CNNs, according to Fig.
@@
 Nonetheless, unlike for CNNs, a recently proposed faithful LRP composite [2] designed for the ViT-B-16 model (detailed in Sec.
-0.A.5), is not ideal for pruning.
+A.5), is not ideal for pruning.
 This finding, and our experiment in Sec.
-0.F.2 show that optimizing an attributor for two different contexts of faithfulness (input or latent space/pruning) (Sec.
+F.2 show that optimizing an attributor for two different contexts of faithfulness (input or latent space/pruning) (Sec.
 3.3) does not necessarily lead to an attributor that attributes faithfully in both input and latent space.
 The poorer performance of heuristically tuned LRP from [48] (Sec.
-0.A.3.3) compared to random when pruning attention heads, underscores both the greater challenge of pruning this structure and the significance of our proposed optimization procedure.
+A.3.3) compared to random when pruning attention heads, underscores both the greater challenge of pruning this structure and the significance of our proposed optimization procedure.
+4.4 How Pruning Affects Model Explanations
+In addition to keeping track of the model performance, the field of XAI encourages us in the following to investigate the model behavior by observing explanation heatmaps.
+Concretely, we expect in the ideal pruning scheme, that heatmaps change as late as possible when increasing the pruning rate.
+This reflects that the task-relevant components are retained as long as possible.
+As an illustrative example, we prune the attention heads of the ViT-B-16 model pre-trained on ImageNet with the aim to predict ImageNet corgi classes (“Pembroke Welsh” and “Cardigan Welsh”) as shown in Fig.
+7.
+As a quantitative measure, we compute the cosine similarity between the original heatmap (using the recently proposed composite of [1]) and the heatmap of the pruned model for different pruning rates over the validation set.
 Figure 7: Pruning a ViT with the aim to retain a high accuracy for detecting ImageNet corgi classes (“Pembroke Welsh” and “Cardigan Welsh”).
@@
 Surprisingly, the unoptimized LRP composite performs even worse than random pruning, stressing the need to optimize attribution methods before applying them heuristically.
-4.4 How Pruning Affects Model Explanations
-In addition to keeping track of the model performance, the field of XAI encourages us in the following to investigate the model behavior by observing explanation heatmaps.
-Concretely, we expect in the ideal pruning scheme, that heatmaps change as late as possible when increasing the pruning rate.
-This reflects that the task-relevant components are retained as long as possible.
-As an illustrative example, we prune the attention heads of the ViT-B-16 model pre-trained on ImageNet with the aim to predict ImageNet corgi classes (“Pembroke Welsh” and “Cardigan Welsh”) as shown in Fig.
-7.
-As a quantitative measure, we compute the cosine similarity between the original heatmap (using the recently proposed composite of [1]) and the heatmap of the pruned model for different pruning rates over the validation set.
 On the one hand, we can see that random pruning or unoptimized attribution-based pruning leads to a much earlier change in heatmaps compared to our optimized LRP-based approach.
@@
 Overall, using local XAI methods for pruning irrelevant model components demonstrates high potential in our experiments.
+Acknowledgements
+This work was supported by the Federal Ministry of Education and Research (BMBF) as grant BIFOLD (01IS18025A, 01IS180371I); the German Research Foundation (DFG) as research unit DeSBi (KI-FOR 5363); the European Union’s Horizon Europe research and innovation programme (EU Horizon Europe) as grant TEMA (101093003); the European Union’s Horizon 2020 research and innovation programme (EU Horizon 2020) as grant iToBoS (965221).
 Appendix
-Appendix 0.A Attribution Methods
+Appendix A Attribution Methods
 As it has been discussed in Sec.
@@
 We focus mostly on LRP and it will be compared with Integrated Gradient as a commonly used gradient-based explainer.
-0.A.1 Integrated Gradient
+A.1 Integrated Gradient
 As an improvement to pure gradient, Integrated Gradient [43] interpolates input $x$ for $m$ steps and computes their gradients sequentially.
@@
 For the experiments conducted in this paper, we interpolate each input for $20$ steps.
-0.A.2 Variants of LRP
+A.2 Variants of LRP
 In addition to the vanilla rule of LRP demonstrated in 6, [4, 32] later on proposed different rules as an extension of vanilla rule designed to serve different purposes.
 Most common extension of LRP will be discussed based on the layer type they target.
-0.A.3 Tackling Linear and Convolution Layers
+A.3 Tackling Linear and Convolution Layers
 As a side note, in the propagation process of LRP, linear layers of DNNs such as fully connected, and convolution layers, are treated similarly.
-0.A.3.1 LRP-$\epsilon$
+A.3.1 LRP-$\epsilon$
 The most fundamental problem of rule 6 causing computational instability, is division by zero which takes place when a neuron is not activated at all ($z_{j}=0$).
@@
 As a side note, sign(0) = 1. Although the value of $\epsilon$ can be tuned, we set it to $61e-$ 6 in every use-case.
-0.A.3.2 LRP-$\alpha\beta$
+A.3.2 LRP-$\alpha\beta$
 As described in figure Fig.
@@
 $\centering R_{i\leftarrow j}=\bigr{(}\alpha\frac{z^{+}_{ij}}{z^{+}_{j}}+\beta\frac{z^{-}_{ij}}{z^{-}_{j}}\bigr{)}R_{j}\@add@centering$ (A.4)
-0.A.3.3 LRP-$z^{+}$
+A.3.3 LRP-$z^{+}$
 An extreme case for the above formulation when it is required to ignore all negative relevance, is the combination of $\alpha=1$ and $\beta=0$, which is referred to as LRP-$z^{+}$ used by [48].
-0.A.3.4 LRP-$\gamma$
+A.3.4 LRP-$\gamma$
 Another derivation, with the same purpose of LRP-$\alpha\beta$ to control over the distribution of positive and negative relevance separately, has this form:
@@
 However, to prevent further computational costs unlike results from [2] in Sec.
-0.A.5, we refrain from tuning it.
-0.A.4 Tackling Attention and Softmax Non-linearity
+A.5, we refrain from tuning it.
+A.4 Tackling Non-linearities
+A.4.1 Activation Functions
+In LRP framework, a non-linear activation function $\sigma$ (e.g., ReLU), with $\sigma_{i}(z_{i})=a_{i}$ is handled similar to Eq.
+6, where $z_{ij}=\delta_{ij}z_{i}$ and $\delta$ represents the Kronecker delta.
+A.4.2 Attention and Softmax on Transformers
 The key component of the attention module [44], consists of two equations:
@@
 $b$ is the batch dimension including the number of heads, and $d_{k},d_{v}$ indicate the embedding dimensions, and $s_{q},s_{k}$ are the number of query and key/value tokens.
-0.A.4.1 CP-LRP
+A.4.3 CP-LRP
 In the extension of LRP from [3], it was proposed to regard the attention matrix (softmax output) A in Eq.
@@
 A.3.
-0.A.4.2 AttnLRP
+A.4.4 AttnLRP
 The recent work of [2] extends LRP to Large Language Models and Vision Transformers by proposing to linearize the softmax function in Eq.
@@
 where $J$ is the Jacobian of the softmax computed at the current point $x$ and $\tilde{b}$ is the bias term of the linearization. For more details, please refer to [2].
-0.A.5 Faithful LRP Composites
+A.5 Faithful LRP Composites
 The works of [24, 2] introduced 2 different LRP composites as faithful explainer [37].
@@
 $\gamma$ $\gamma$ $\epsilon$ $\epsilon$ Attn-LRP$(z^{+})$
-Appendix 0.B Spatial Dimension of To-Be-Pruned Components
+Appendix B Spatial Dimension of To-Be-Pruned Components
 In Eq.
@@
 $\centering\bar{R_{p}}=\sum_{i}^{n}\sum_{j}^{d_{q}}\big{|}\sum_{l}^{d_{k}}R_{(p,j,l)}({x_{i}})\big{|}\@add@centering$ (A.14)
-Appendix 0.C Additionally on Overparameterization
+Appendix C Additionally on Overparameterization
 Similar to Fig.
@@
 Additionally, Relevance denoted in the legend is computed by LRP framework with $\epsilon$ rule (see Sec.
-0.A.2) applied across all layers.
-Appendix 0.D On Optimization: From Bayesian to Grid Search
+A.2) applied across all layers.
+Appendix D On Optimization: From Bayesian to Grid Search
 To find an optimal LRP composite for pruning, our approach takes place by discovering prospective parameters (Sec.
@@
 This seems to be a more effective solution rather than naively applying Grid Search over the whole parameters when we can reduce from $50\%$ up to $90\%$ of our search space approximately.
-Appendix 0.E On Number of Reference Samples
+Appendix E On Number of Reference Samples
 Fig.
@@
 We used LRP-$\epsilon$ over all layers.
-Appendix 0.F Optimized Composites
-0.F.1 Top Composites for Pruning
+Appendix F Optimized Composites
+F.1 Top Composites for Pruning
 In later tables, based on the conducted optimization, we demonstrate top composites for pruning on CNNs (Tab.
@@
 ViT-B-16 (Attention Heads) 0.012 0.048 0.025 0.029 0.012 0.061
-0.F.2 Evaluating the Composite
+F.2 Evaluating the Composite
 As described in Sec.
@@
 4.3 exhibited that LRP composite (Sec.
-0.A.5) suggested by [2] is not the most suitable option for pruning.
+A.5) suggested by [2] is not the most suitable option for pruning.
 However, qualitative and quantitative experiment (based on [5, 4, 2]) in Fig.
@@
 This experiment took place by using around $1000$ samples from ImageNet and $100$ perturbation steps (see work of [7] for the description of this metric).
-Appendix 0.G Layer-wise Relevance Flow
+Appendix G Layer-wise Relevance Flow
 Later figures (A.4, A.5, A.6, A.7, A.8, A.9) partially elucidate the pruning strategy by illustrating relevance magnitude of components in a layer-wise fashion.
@@
 Figure A.9: Similar to Fig. A.8 there are in total, 4224 convolution filters in VGG-16-BN distributed in 13 layers. Higher relevance magnitude is visible at mid to lower-level convolution filters.
-Appendix 0.H Change in Behaviour
+Appendix H Change in Behaviour
 In later figures (A.10, A.11, A.12), we qualitatively demonstrate examples of pruning based on 3 randomly chosen classes and trace changes in the explanation heatmap of the pruned model.
 Figure A.10: Changes in ResNet-18 decision-making for domain restriction of 3 randomly chosen classes (Box Turtle, Alligator Lizard, and Moving Van) via LRP heatmaps (using composite described in Sec.
-0.A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and background despite considering the main object itself.
+A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and background despite considering the main object itself.
 Heatmaps of the model with high pruning rates demonstrate an effect caused by a subsampling shortcut connection becoming apparent.
 Figure A.11: Changes in ResNet-50 decision-making for domain restriction of 3 randomly chosen classes (Flamingo, Gordon Setter, and Snowplow) via LRP heatmaps (using composite described in Sec.
-0.A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and unlike ResNet-18 ignores the main object.
+A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and unlike ResNet-18 ignores the main object.
 The shortcut connections of ResNet architecture result in checkerboard pattern exhibited in the heatmaps.
 Figure A.12: Changes in VGG-16-BN decision-making for domain restriction of 3 randomly chosen classes (Digital Clock, Patio, Terrace, and Prayer Rug) via LRP heatmaps (using composite described in Sec.
-0.A.5) indicate lower pruning capabilities of this architecture compared to ResNet-18 and ResNet-50.
+A.5) indicate lower pruning capabilities of this architecture compared to ResNet-18 and ResNet-50.